OpenStudy (anonymous):

What is the vertex of the graph of (x-3)^2=y+4?

OpenStudy (anonymous):

Is it (3, -4)???

OpenStudy (tkhunny):

Eyeball problem. Just look at it and write it down. That's why we call it the "Vertex Form". Listen to ILC.

OpenStudy (anonymous):

.... I am asking for the vertex. The reason why I posted this is because I don't know how exactly to find it. If I knew how to find the vertex I would use my eyeballs to find it in the equation. @tkhunny

OpenStudy (tkhunny):

?? Very strange reply. You looked at it and wrote down the Vertex. I told you to trust your eyes. What question do you still have? Always: (x-h) and (y-k) are in there in the Vertex Form. The Vertex is (h,k). (x-3)^2=y+4? x-3 leads immediately to h = 3 y+4 takes a moment to conclude the k = -4 Eyeball problem: (3,-4)